# Rational motives on pro-algebraic stacks

### v3

@@ 2. $\bullet$
One might wonder why we chose $\textup{DM}^{*}$ and not $\textup{DM}^{!}$. Our main reason is convenience. For monoidal $-{\mathcal{I}}$-algebraic stacks, we conjecture that we have an equivalence between both possible definitions. But by working with $\textup{DM}^{*}$, monoidality of $\textup{DM}(x,X)$ follows straightforwardly. We want to remark that with this definition of $\textup{DM}(x,X)$, we recover the usual $\textup{DM}(X)$ if $(x,X)$ is classical monoidal.
In rest of this section is we want to prove Theorem 5.
@@ 2.1 Pullback formalisms
We will first define pullback formalisms on schemes after Drew-Gallauer (cf. [DG22]). Afterward, we want to show that DM on $-{\mathcal{I}}$-algebraic stacks defines a pullback formalism. We will analyze such structures more generally in this subsection without using $-{\mathcal{I}}$-algebraic stacks.
@@
Let ${\mathcal{E}}$ be an $\infty$-category that admits finite limits and let us fix a class of morphisms $P\subseteq\operatorname{Fun}(\Delta^{1},{\mathcal{E}})$ that is closed under pullback and equivalences and composition. A pullback formalism on ${\mathcal{E}}$ with respect to $P$ is a functor
${\mathcal{M}}^{*,\otimes}\colon{\mathcal{E}}^{\operatorname{op}}\rightarrow\textup{CAlg}(\textup{Pr}^{L,\otimes}),\quad f\mapsto f^{*}\quad\textup{(we denote the right adjoint of $f^{*}$ by $f_{*}$)}$
@@
    If a morphism $f\colon X\rightarrow Y$ of $S$ lies in $P$, then there exists a left adjoint $f_{\sharp}$ of $f^{*}$.
@@
    ${X\times_{Z}Y}$${X}$${Y}$${Z}$$\scriptstyle{g}$$\scriptstyle{f}$$\scriptstyle{f}$$\scriptstyle{g}$
    in ${\mathcal{S}}$ such that $f$ is in $P$, the exchange transformation
@@
    $f_{\sharp}(M\otimes_{{\mathcal{M}}^{*,\otimes}(X)}f^{*}N)\rightarrow f_{\sharp}M\otimes_{{\mathcal{M}}^{*,\otimes}(Y)}N$
    is an equivalence, for any $M\in{\mathcal{M}}^{*,\otimes}(X)$ and $N\in{\mathcal{M}}^{*,\otimes}(Y)$.
Let us remark that this definition of a pullback formalism differs slightly from the definition of op. cit.. This adjustment is necessary for our means and a for a more general overview we refer to op. cit..
@@
For property (2) in Definition 2.9 we may equivalently ask for the *-exchange morphism $f^{*}g_{*}\rightarrow g^{\prime}_{*}f^{\prime*}$ to be an equivalence (cf. [Ayo07, Prop. 1.1.9]).
@@
Throughout this article, we will be mainly interested in the case where ${\mathcal{E}}\simeq\textup{Stk}^{\textup{lft}}_{S}$, where $\textup{Stk}^{\textup{lft}}_{S}$ denotes the category of locally of finite type Artin stacks over $S$ (recall $S$ from our notations), $P$ the class of smooth morphisms and ${\mathcal{M}}^{*,\otimes}=\textup{DM}_{\mathbb{Q}}$ with the *-pullback. If ${\mathcal{E}}$ is the category of finite type $S$-schemes, instead of $\textup{DM}_{\mathbb{Q}}$, we could also work $\textup{DM}_{\mathbb{Z}}$ or even with the stable homotopy category SH.
Definition 2.12.
A pullback formalism ${\mathcal{M}}^{*,\otimes}$ is called compactly generated if ${\mathcal{M}}^{*,\otimes}$ factors through the functor $\textup{CAlg}(\textup{Pr}^{L,\otimes}_{\omega})$, where $\omega$ denotes the cardinality of the natural numbers.
Example 2.13.
Recall that in the assumptions 1.1, we fixed a regular scheme $S$ of finite type over a excellent noetherian scheme $B$ of dimension $\leq 2$. By [CD19, Thm. 16.1.4, Cor. 6.2.2] the $\infty$-category $\textup{DM}(X,\mathbb{Q})$ is compactly generated for any finite type $S$-scheme and $f^{*}$ preserves compact objects. In particular, following Example 2.11, $\textup{DM}_{\mathbb{Q}}$ induces a compactly generated pullback formalism.
Notation 2.14.
For this subsection, we fix an $\infty$-category ${\mathcal{E}}$ that admits finite limits, a class of morphisms $P_{{\mathcal{E}}}$ that is stable under pullbacks, compositions and equivalences. Further, we fix a compactly generated pullback formalism ${\mathcal{M}}^{*,\otimes}$ on ${\mathcal{E}}$ with respect to $P_{{\mathcal{E}}}$ and simply denote it by ${\mathcal{M}}^{\otimes}$.
Example 2.15.
Let us give $2$ examples for ${\mathcal{E}}$, $P_{{\mathcal{E}}}$ and ${\mathcal{M}}^{\otimes}$ that satisfy our assumptions. Recall $S$ from our assumptions and notations 1.1.
  1. (1)
    Let ${\mathcal{E}}=\textup{Sch}_{S}^{\textup{lft}}$ be the category of $S$-schemes locally of finite type and $P_{{\mathcal{C}}}$ the class of smooth morphisms in $\textup{Sch}_{S}$. Then $\textup{DM}_{\mathbb{Z}}$ together with *-pullbacks satisfies our assumptions.
  2. (2)
    Let us consider $\textup{DM}_{\mathbb{Q}}$ together with *-pullbacks. Then $\textup{DM}_{\mathbb{Q}}$ satisfies étale descent and standard arguments show that we can take ${\mathcal{E}}=\textup{Stk}^{\textup{lft}}_{S}$ and $P_{{\mathcal{E}}}$ to be the class of smooth morphisms (cf. [LZ12]).
We will start by extending ${\mathcal{M}}^{\otimes}$ to diagrams in $K$, where $K$ is a simplicial set, via taking colimits. By the remark below $\textup{CAlg}(\textup{Pr}^{L,\otimes}_{\omega})$ admits colimits.
@@
Let us remark that $\Pr^{L}_{\omega}$ is presentable by [Lur17, Lem. 5.3.2.9], thus so is $\textup{Mod}_{\operatorname{Sp}}(\textup{Pr}^{L,\otimes}_{\omega})$ by [Lur17, Cor. 4.2.3.7] and hence also $\textup{CAlg}(\textup{Pr}^{L,\otimes}_{\omega})$ by [Lur17, Prop. 3.2.3.5].
Let $K$ be a simplicial set. Then ${\mathcal{M}}$ induces a functor
$\operatorname{Fun}(K,{\mathcal{E}}^{\operatorname{op}})\rightarrow\operatorname{Fun}(K,\textup{CAlg}(\textup{Pr}^{L,\otimes}_{\omega})).$
Composing this functor with the colimit functor, we obtain
${\mathcal{M}}^{\otimes}_{K}\colon\operatorname{Fun}(K,{\mathcal{E}}^{\operatorname{op}})\rightarrow\textup{CAlg}(\textup{Pr}^{L,\otimes}_{\omega}).$
If $(x,X)$ is a $-{\mathcal{I}}$-algebraic stack, we see that the following notation agrees Definition 2.8.
Notation 2.17.
Let ${\mathcal{N}}$ be a pullback formalism on locally of finite type Artin $S$-stacks with respect to smooth morphisms. Let $(x,X)$ be a $-{\mathcal{I}}$-algebraic stack. Then we define
${\mathcal{N}}(x,X)\coloneqq{\mathcal{N}}_{{\mathcal{I}}}^{\otimes}(x).$
@@
Let $K$ be a filtered $\infty$-category. Let ${\mathcal{P}}_{K}$ denote the class of morphisms in $\operatorname{Fun}(K,{\mathcal{E}})$ such that $f\in\operatorname{Fun}(K,{\mathcal{E}})$ belongs to ${\mathcal{P}}_{K}$ if and only if each $f_{k}$ belongs to $P_{{\mathcal{E}}}$ and each square in ${\mathcal{M}}_{K}(f)$ is right adjointable, i.e. for any morphism $i\rightarrow j\in K$ the induced morphism $f_{i}^{*}x_{ij*}\rightarrow y_{ij*}f_{j}^{*}$ is an equivalence (cf. [Lur17, Def. 4.7.4.13]).
Then ${\mathcal{M}}_{K}$ is a pullback formalism on $\operatorname{Fun}(K,{\mathcal{E}})$ for the class ${\mathcal{P}}_{K}$.
Proposition 2.18 is one of the crucial points in extending the motivic six functor formalism on DM for Artin-stacks to $-{\mathcal{I}}$-algebraic stacks. But before we can prove this proposition, we need some additional lemmas.
@@
Let $K$ a filtered $\infty$-category. Let $F\colon{\mathcal{E}}\rightarrow\Pr^{L}_{\omega}$ be a functor. Denote by $F_{K}$ the composition of $\operatorname{Fun}(K,F)$ with the colimit functor. Let $X$ be in $\operatorname{Fun}(E,{\mathcal{C}})$. Then we have
$F_{K}(X)\simeq\textup{Ind}(\mathop{\textup{colim}}_{i\in K,x^{*}}F(X_{i})^{c})$
and any $M\in F_{K}(X)$ is compact if and only if there exists an $i\in K$ and $M_{i}\in F(X_{i})^{c}$ that lifts $M$.
@@
The first assertion follows from (cf. [GR17, Cor. 7.2.7]), where the colimit on the right is taken in $\textup{Cat}_{\infty}$.
Note that the filtered colimit of idempotent complete $\infty$-categories is idempotent complete (cf. [Lur09, Cor. 4.4.5.21]). Thus, any compact object comes from an object of the colimit on the right hand side (cf. [Lur09, Lem. 5.4.2.4] and [Lur21, 02LG]). ∎
@@
Let $K$ be a $\infty$-category. Let $F\colon{\mathcal{E}}\rightarrow\Pr^{L}_{\omega}$ be a functor. Denote by $F_{K}$ the composition of $\operatorname{Fun}(K,F)$ with the colimit functor. Then we have the following.
  1. (1)
    Let $X\in\operatorname{Fun}(K,{\mathcal{E}})$. We denote by $X^{\operatorname{op}}$ the opposite functor. Further, let us denote by $F_{*}\colon{\mathcal{E}}^{\operatorname{op}}\rightarrow\Pr^{R}$ the functor induced by the equivalence $\Pr^{L}\simeq(\Pr^{R})^{\operatorname{op}}$ (cf. [Lur09, Cor. 5.5.3.4]). Then we have
    $F_{K}(X)\simeq\lim_{\Pr^{R}}F_{*}\circ X^{\operatorname{op}},$
    where the equivalence is in $\textup{Cat}_{\infty}$. More precisely, $F_{K}$ is equivalent to the right Kan extension of $F_{*}$ along $\delta^{\operatorname{op}}$.
  2. (2)
    If $K$ be a filtered $\infty$-category, then for any $k_{0},k_{1}\in K$ the composition
    $F(X_{k_{0}})\xrightarrow{\textup{ins}_{k_{0}}^{X}}F_{K}(X)\simeq\lim_{\Pr^{R}}F_{*}\circ X\xrightarrow{p_{k_{1}}^{X}}{\mathcal{M}}(X_{k_{1}})$
    is canonically equivalent to
    $\mathop{\textup{colim}}_{k\in K,\alpha\colon k_{0}\rightarrow k,\beta\colon k_{1}\rightarrow k}F_{*}(X_{\beta})\circ F_{K}(X_{\alpha})$
    where the colimit is taken in $\operatorname{Fun}_{\Pr^{L}}(F(X_{k_{0}}),F(X_{k_{1}}))$.
Proof.
Remark 2.21.
The equivalence in Lemma 2.20 for $X\in\operatorname{Fun}(K,{\mathcal{E}})$ is induced via the following process (cf. [Gai12, Lem. 1.3.2]). The projections $p^{X}_{k}\colon\lim_{\Pr^{R}}F_{*}\circ X\rightarrow X(k)$ admit left adjoint, denoted by $\gamma_{k}^{X}$. These left adjoints assemble to an equivalence
(2.21.1) $h_{X}\colon F_{K}(X)\xrightarrow{\sim}\lim_{k\in K^{\operatorname{op}},x_{*}}F(X_{i}).$
This is an equivalence in $\Pr^{L}$, where we view the RHS naturally as an object in $\Pr^{L}$. The inverse of $h_{X}$ is equivalent to $\mathop{\textup{colim}}_{k\in K}\textup{ins}^{X}_{k}\circ p_{k}^{X},$ where $\textup{ins}^{X}_{k}\colon F(X_{k})\rightarrow{\mathcal{M}}_{K}(X)$ denotes the canonical map.
Lemma 2.22.
Consider the setting of Lemma 2.20 (2). If $f\colon X\rightarrow Y$ is a morphism in $\operatorname{Fun}(K,{\mathcal{E}})$ such that each $f_{k}$ admits a left adjoint for all $k\in K$, then $f^{*}\coloneqq F(f)$ admits a left adjoint.
@@
By Lemma 2.20 (1), we see that $f_{*}\colon\lim_{K}F_{*}\circ X\rightarrow\lim_{K}F_{*}\circ Y$ admits a left adjoint $f^{*}$, that has to be compatible with the identification of $\lim_{K}F_{*}$ and $F_{K}$. We claim that the following diagram is commutative (up to homotopy)
${\lim_{K}F_{*}\circ Y}$${F(Y_{i})}$${\lim_{K}F_{*}\circ X}$${F(X)}$$\scriptstyle{p_{i}^{Y}}$$\scriptstyle{f^{*}}$$\scriptstyle{f_{i}^{*}}$$\scriptstyle{p_{i}^{X}}$
for all $i\in K$. This claim implies that $f^{*}$ commutes with limits and therefore admits a left adjoint $f_{\sharp}$.

### v4

@@ 3. $\bullet$
One might wonder why we chose $\textup{DM}^{*}$ and not $\textup{DM}^{!}$. Our main reason is convenience. We conjecture that we have an equivalence between both possible definitions. But by working with $\textup{DM}^{*}$, monoidality of $\textup{DM}(x,X)$ follows by definition. Also, we will see that if $\mathcal{I}$$\simeq\mathbb{N}$, then both possible definitions are equivalent.
Let us now formalize Theorem 3 of the introduction. We will postpone the technical proof to Section 3.
Theorem 2.9.
Let $f\colon(x,X)\rightarrow(y,Y)$ be a morphism of pro-$\mathcal{I}$}-algebraic stacks. Then there is an adjunction
$f^{*}\colon\textup{DM}(y,Y)$$\textup{DM}(x,X)\colon f_{*},$}
where $f^{*}$ is symmetric monoidal and preserves compact objects. The functor $f^{*}$ extends to a functor $(x,X)\mapsto\textup{DM}(x,X)$. Moreover, the functor DM satisfies the following property on pro-$\mathcal{I}$}-algebraic stacks.
  1. (PB)
    If $f$ is smooth and *-adjointable, then $f^{*}$ admits a left adjoint $f_{\sharp}$ satisfying base change and projection formula.
Assume now that $f$ locally of finite type. Further, assume one of the following holds true
  1. ($\ast$)
    $f$ is cartesian, or
  2. ($\dagger$)
    $\mathcal{I}$$\simeq\mathbb{N}$, $f$ is !!-adjointable, $(x,X)$ and $(y,Y)$ are tame.
Then we can upgrade $f^{*}$ to be part of a $6$-functor formalism in the following sense. There exists an adjunction
$f_{!}\colon\textup{DM}(x,X)$$\textup{DM}(y,Y)\colon f^{!},$}
where $f_{!}$ defines a covariant functor on pro-algebraic stacks satisfying ()(*) or $(\dagger)$. Further, the adjunctions above satisfy the following properties.
  1. (F1)
    If $f$ is cartesian and proper, we have a canonical equivalence $f_{*}\simeq f_{!}$.
  2. (F2)
    The functor $f^{!}$ commutes with arbitrary colimits.
  3. (F3)
    (Base change) For any pullback square
[…]
    of pro-algebraic stacks with morphisms satisfying ()(*) resp. $(\dagger)$, the exchange morphisms
    $f^{*}g_{!}\rightarrow g^{\prime}_{!}f^{\prime*}$
    $g^{!}f_{*}\rightarrow f^{\prime}_{*}g^{\prime!}$
    are equivalences.
  4. (F4)
Furthermore, the following properties hold.
  1. (M1)
    (Homotopy invariance) Let us define $(q,\mathbb{A}^{1}_{X})$ via the system $i\mapsto\mathbb{A}^{1}_{X_{i}}$. Then $p^{*}$ is fully faithful, where $p\colon(q,\mathbb{A}^{1}_{X})\rightarrow(x,X)$ denotes the projection.
  2. (M2)
    $i_{!}i^{!}\rightarrow\operatorname{id}\rightarrow j_{*}j^{*}$
    $j_{!}j^{!}\rightarrow\operatorname{id}\rightarrow i_{*}i^{*}.$
Assuming that $\mathcal{I}$} admits an initial object $0\in{\mathcal{I}}$, we moreover have the following properties.
  1. (M3)
  2. (M4)
  3. (M5)
    (Purity) If $f$ is cartesian, we have a canonical equivalence $f^{!}\simeq\operatorname{\textup{Th}}(\Omega_{f_{0}})1_{X}\otimes f^{*}$.
If we further restrict ourselves to pro-$\mathcal{I}$}-schemes, i.e. inverse systems of schemes locally of finite type over $S$ indexed over $\mathcal{I}$}, then $f^{*}$ extends to a $6$-functor formalism in the sense of Liu-Zheng [LZ17].
Before we continue with the proof, we want to show that $\textup{DM}(x,X)\simeq\textup{DM}(X)$ for classical pro-algebraic stacks, as one would expect. We will need the following result.
Lemma 2.10.
Let $\mathcal{E}$} be an $\infty$-category that admits finite limits. Let $F\colon{\mathcal{E}}\rightarrow\Pr^{L}_{\omega}$ be a functor. Denote by $F_{{\mathcal{I}}}$ the composition of $\operatorname{Fun}({\mathcal{I}},F)$ with the colimit functor $\operatorname{Fun}({\mathcal{I}},\Pr^{L}_{\omega})\rightarrow\Pr^{L}_{\omega}$. Then we have the following.
  1. (1)
    Let $X\in\operatorname{Fun}({\mathcal{I}},{\mathcal{E}})$. We denote by $X^{\operatorname{op}}$ the opposite functor. Further, let us denote by $F_{*}\colon{\mathcal{E}}^{\operatorname{op}}\rightarrow\Pr^{R}$ the functor induced by the equivalence $\Pr^{L}\simeq(\Pr^{R})^{\operatorname{op}}$ [Lur09, Cor. 5.5.3.4]. Then we have
    $F_{{\mathcal{I}}}(X)\simeq\lim_{\Pr^{R}}F_{*}\circ X^{\operatorname{op}},$
    where the equivalence is in $\textup{Cat}_{\infty}$.
  2. (2)
    For any $i_{0},i_{1}\in{\mathcal{I}}$ the composition
    $F(X_{i_{0}})\xrightarrow{\textup{ins}_{i_{0}}^{X}}F_{{\mathcal{I}}}(X)\simeq\lim_{\Pr^{R}}F_{*}\circ X\xrightarrow{p_{i_{1}}^{X}}F(X_{i_{1}})$
    is canonically equivalent to
    $\mathop{\textup{colim}}_{k\in{\mathcal{I}},\alpha\colon i_{0}\rightarrow k,\beta\colon i_{1}\rightarrow k}F_{*}(X_{\beta})\circ F_{{\mathcal{I}}}(X_{\alpha})$
    where the colimit is taken in $\operatorname{Fun}_{\Pr^{L}}(F(X_{i_{0}}),F(X_{i_{1}}))$.
Proof.
Remark 2.11.
The equivalence in Lemma 2.10 for $X\in\operatorname{Fun}({\mathcal{I}},{\mathcal{E}})$ is induced via the following process [Gai12, Lem. 1.3.2]. The projections $p^{X}_{i}\colon\lim_{\Pr^{R}}F_{*}\circ X\rightarrow X_{i}$ admit left adjoints, denoted by $\gamma_{i}^{X}$. These left adjoints assemble to an equivalence
(2.11.1) $h_{X}\colon F_{{\mathcal{I}}}(X)\xrightarrow{\sim}\lim_{i\in{\mathcal{I}}^{\operatorname{op}},x_{*}}F(X_{i}).$
This is an equivalence in $\Pr^{L}$, where we view the RHS naturally as an object in $\Pr^{L}$. The inverse of $h_{X}$ is equivalent to $\mathop{\textup{colim}}_{i\in{\mathcal{I}}}\textup{ins}^{X}_{i}\circ p_{i}^{X},$ where $\textup{ins}^{X}_{i}\colon F(X_{i})\rightarrow F_{{\mathcal{I}}}(X)$ denotes the canonical map.
Let us fix the above notation, as we will need it later on.
Notation 2.12.
We fix the notation as in Remark 2.11, i.e. for a pro-$\mathcal{I}$}-algebraic stack $(x,X)$, we denote by $h_{X}$ the equivalence
we denote for any $i\in{\mathcal{I}}$ by
the canonical inclusion resp. projections and we set $\gamma_{i}^{X}\coloneqq h_{X}\circ\textup{ins}_{i}^{X}$.
Proposition 2.13.
Assume that $\mathcal{I}$} admits an initial object $0\in{\mathcal{I}}$. Let $(x,X)$ be a classical pro-$\mathcal{I}$}-algebraic stack. Then the natural map
$\textup{DM}(x,X)\rightarrow\textup{DM}(X)$
is an equivalence of DG-categories.
Proof.
For the proof we will use two immediate facts. The canonical map $\textup{DM}(x,X)\rightarrow\textup{DM}(X)$ is by construction monoidal and preserves colimits. Moreover, to show that $\textup{DM}(x,X)\rightarrow\textup{DM}(X)$ is an equivalence, we may work with the right adjoints
$\textup{DM}(X)\rightarrow\lim_{i\in{\mathcal{I}}^{\operatorname{op}},x_{*}}\textup{DM}(X_{i})$
(this follows from Lemma 2.10 applied to $F=\textup{DM}:\textup{Stk}^{\textup{lft},\operatorname{op}}_{S}\rightarrow\textup{Pr}^{L}_{\omega}$). By assumption $X_{0}$ admits a smooth cover by affine finite type $S$-schemes $U_{0,k}=\operatorname{Spec}(A_{k})$, for some set $K$ and $k\in K$. As $\textbullet X_{i}\rightarrow X_{0}$ is affine for all $i\in{\mathcal{I}}$, the pullback $U_{i,k}\coloneqq X_{i}\times_{X_{0}}U_{i,k}$ is also a smooth cover by affines. The same holds after passage to limits, i.e. for $U_{k}\coloneqq X\times_{X_{0}}U_{0,i}$ the projection $\coprod_{k\in K}U_{i}\rightarrow X$ is an effective epimorphism [Sta25, 01YZ]. In particular, as DM satisfies $h$-descent (cf. [RS20, Thm. 2.2.16]), we have $\lim_{\Delta}\textup{DM}(\check{C}_{\bullet}(\coprod_{k\in K}U_{i}/X))\simeq\textup{DM}(X)$. By construction, each of the $U_{i}$ is equivalent to $\lim_{i\in{\mathcal{I}}}U_{i,k}$. As the $U_{i,k}$ are affine of finite type over $S$, we have
$\textup{DM}(U_{i})\simeq\textup{DM}(\lim_{i\in{\mathcal{I}}}U_{i,k})\simeq\mathop{\textup{colim}}_{i\in{\mathcal{I}},x^{*}}\textup{DM}(U_{i,k})\simeq\lim_{i\in{\mathcal{I}}^{\operatorname{op}},x_{*}}\textup{DM}(U_{i,k}).$
This yields the equivalence
$\textup{DM}(\check{C}_{\bullet}(\coprod_{k\in K}U_{i}/X))\simeq\lim_{i\in{\mathcal{I}},x_{*}}\textup{DM}(\check{C}_{\bullet}(\coprod_{k\in K}U_{i,k}X_{i})).$
Putting all of this together, we claim that this yields
$\textup{DM}(X)\simeq\lim_{\Delta}\textup{DM}(\check{C}_{\bullet}(\coprod_{k\in K}U_{i}/X))$ $\simeq\lim_{i\in{\mathcal{I}}^{\operatorname{op}},x_{*}}\lim_{\Delta}\textup{DM}(\check{C}_{\bullet}(\coprod_{k\in K}U_{i,k}/X_{i}))$
as desired.
$\textup{DM}(\check{C}_{\bullet}(\coprod_{k\in K}U_{j,k}/X_{j}))\rightarrow\textup{DM}(\check{C}_{\bullet}(\coprod_{k\in K}U_{i,k}/X_{i})).$
But this follows immediately from smooth base change, as
$\check{C}_{\bullet}(\coprod_{k\in K}U_{j,k}/X_{j})\simeq\check{C}_{\bullet}(\coprod_{k\in K}U_{i,k}/X_{i})\times_{X_{i}}X_{j}.$
Remark 2.14.
Let $Y$ be a scheme and let $(x,X)$ be a classical pro-algebraic stack such that each $X_{i}$ is representable by a finite type $Y$-scheme. The proof of Proposition 2.13 can be repeated to obtain the following more general result on schemes.
  1. $\bullet$
    Let $F\colon\textup{Sch}^{\textup{ft},\operatorname{op}}_{Y}\rightarrow\textup{CALg}(\textup{Pr}^{L,\otimes})$ be a Zariski sheaf, then $\mathop{\textup{colim}}F\circ x\simeq F(X)$.
This includes for example motives with integral coefficients and the stable homotopy category of Morel-Voevodsky.
By definition if $(x,X)$ is a strict pro-algebraic stack then the transition maps $x_{ij}^{*}$ are fully faithful for all $i\rightarrow j\in{\mathcal{I}}$. In particular, we would expect that the pullback $\textup{DM}(X_{0})\rightarrow\textup{DM}(X)$ is fully faithful. At least when we work with $\textup{DM}(x,X)$ this is true, as our explicit analysis will show. If $(x,X)$ is classical, we see in particular that $\textup{DM}(X_{0})\rightarrow\textup{DM}(X)$ is indeed fully faithful.
Remark 2.15 (Underling motive of strict pro-algebraic stacks).
Let us show how to compute the underlying motive in $\textup{DM}(S)$ of a strict pro-$\mathcal{I}$}-algebraic stack, when $\mathcal{I}$} admits an initial object 0.
Let $(x,X)$ be a strict pro-algebraic stack. Then by Example 2.6 (2), the projection $c_{0}\colon(x,X)\rightarrow X_{0}$ is an adjointable map of pro-algebraic stacks. Thus, by Theorem 2.9, we obtain an adjunction
$c_{0*}\colon\textup{DM}(x,X)$$\textup{DM}(X_{0})\colon c_{0}^{*}$}
Moreover, we see from explicit computations in Lemma 2.10 (2) that $c_{0}^{*}$ is fully faithful, since $c_{0}^{*}$ corresponds to canonical map $\textup{DM}(X_{0})\rightarrow\textup{DM}(x,X)$ and $c_{0*}$ to the projection.
In particular, we will see ath if $f\colon(x,X)\rightarrow S$ is strict, tame and $f_{0}\colon X_{0}\rightarrow S$ denote the structure map, we have for any $M\in\textup{DM}(S)$ the equivalence $f_{!}f^{!}M\simeq f_{0!}f_{0}^{!}M$.
3 (Motivic) six functor formalism on pro-algebraic stacks
In this section we will prove Theorem 2.9.
Moreover, we will not restrict us to the category of algebraic stacks but rather work in the most general context, as we will only need formal properties, such as a geometric setup in the sense of Mann [Man22]. While this is a more abstract formalism we develop, the benefit is that we can also apply this formalism to other settings, for example ind-schemes. Such inverse limits of ind-schemes and étale cohomology were considered by Bouthier-Kazhdan-Varshavsky [BKV22] where they were called placid.
@@ 3.1 Pullback formalisms
We start by analyzing filtered colimits of pullback formalisms after Drew-Gallauer [DG22]. We will analyze such structures more generally in this subsection without using pro-algebraic stacks but work with general diagrams in an $\infty$-category $\mathcal{E}$} admitting finite limits.
@@
Let $\mathcal{E}$} be an $\infty$-category that admits finite limits and let us fix a class of morphisms $P\subseteq\operatorname{Fun}(\Delta^{1},{\mathcal{E}})$ that is closed under pullback, equivalences and composition. A compactly generated pullback formalism on $\mathcal{E}$} with respect to $P$ is a functor
[…]
@@
    If a morphism $f\colon X\rightarrow Y$ of $\mathcal{E}$} lies in $P$, then there exists a left adjoint $f_{\sharp}$ of $f^{*}$.
@@
[…]
    in $\mathcal{E}$} such that $f$ lies in $P$, the exchange transformation
@@
    $f_{\sharp}(M\otimes_{{\mathcal{M}}^{\otimes}(X)}f^{*}N)\rightarrow f_{\sharp}M\otimes_{{\mathcal{M}}^{\otimes}(Y)}N$
    is an equivalence, for any $M\in{\mathcal{M}}^{\otimes}(X)$ and $N\in{\mathcal{M}}^{\otimes}(Y)$.
We call the triple $({\mathcal{M}}^{\otimes},{\mathcal{E}},{\mathcal{P}})$ as above a pullback formalism.
@@
For property (2) in Definition 3.1 we may equivalently ask for the *-exchange morphism $f^{*}g_{*}\rightarrow g^{\prime}_{*}f^{\prime*}$ to be an equivalence [Ayo07, Prop. 1.1.9].
@@
Recall that in the assumptions 1.1, we fixed an excellent noetherian scheme $S$ finite dimension. By [CD19, Thm. 16.1.4, Cor. 6.2.2] the $\infty$-category $\textup{DM}(X,\mathbb{Q})$ is compactly generated for any finite type $S$-scheme and $f^{*}$ preserves compact objects. In particular, $\textup{DM}_{\mathbb{Q}}$ induces a pullback formalism on finite type $S$-schemes. By descent this can be extended to locally of finite type algebraic stacks over $S$ [RS20].
We will start by extending $\mathcal{M}$$^{\otimes}$ to diagrams in $\mathcal{I}$}, via taking colimits.
@@
[…]
Notation 3.5.
Consider the functor
$)\operatorname{Fun}({\mathcal{I}},{\mathcal{E}}^{\operatorname{op}})\rightarrow\operatorname{Fun}({\mathcal{I}},\textup{CAlg}(\textup{Pr}^{L,\otimes}_{\omega}$))
induced by postcomposition with $\mathcal{M}$$^{\otimes}$. Composing this functor with the colimit functor, we obtain
[…]
If $(x,X)$ is a pro-algebraic stack, we see that the following notation agrees Definition 2.8.
Notation 3.6.
Let $\mathcal{M}$$^{\otimes}$ be a pullback formalism on locally of finite type Artin $S$-stacks with respect to smooth morphisms. Let $(x,X)$ be a pro-$\mathcal{I}$}-algebraic stack. Then we define
${\mathcal{M}}(x,X)\coloneqq{\mathcal{M}}_{{\mathcal{I}}}^{\otimes}(x).$
@@
Let $({\mathcal{M}}^{\otimes},{\mathcal{E}},{\mathcal{P}})$ be a pullback formalism. Let $\mathcal{P}$$_{{\mathcal{I}}}$ denote the class of morphisms in $\operatorname{Fun}({\mathcal{I}},{\mathcal{E}}^{\operatorname{op}})$ such that $f\in\operatorname{Fun}({\mathcal{I}},{\mathcal{E}}^{\operatorname{op}})$ belongs to $\mathcal{P}$$_{{\mathcal{I}}}$ if and only if each $f_{i}$ belongs to $\mathcal{P}$} and each square in ${\mathcal{M}}_{{\mathcal{I}}}(f)$ is adjointable, i.e. for any morphism $i\rightarrow j\in{\mathcal{I}}$ the induced morphism $f_{i}^{*}y_{ij*}\rightarrow x_{ij*}f_{j}^{*}$ is an equivalence [Lur17, Def. 4.7.4.13]. Then $\mathcal{M}$$_{{\mathcal{I}}}$ is a pullback formalism on $\operatorname{Fun}({\mathcal{I}},{\mathcal{E}}^{\operatorname{op}})$ for the class $\mathcal{P}$$_{{\mathcal{I}}}$.
Proposition 3.7 is one of the crucial points in extending the motivic six functor formalism on DM for Artin-stacks to pro-algebraic stacks. But before we can prove this proposition, we need some additional lemmas.
@@
Let $F\colon{\mathcal{E}}^{\operatorname{op}}\rightarrow\Pr^{L}_{\omega}$ be a functor. Denote by $F_{{\mathcal{I}}}$ the composition of $\operatorname{Fun}({\mathcal{I}},F)$ with the colimit functor. Let $X$ be in $\operatorname{Fun}({\mathcal{I}},{\mathcal{E}}^{\operatorname{op}})$. Then we have
$F_{{\mathcal{I}}}(X)\simeq\textup{Ind}(\mathop{\textup{colim}}_{i\in{\mathcal{I}},x^{*}}F(X_{i})^{c})$
and any $M\in F_{{\mathcal{I}}}(X)$ is compact if and only if there exists an $i\in{\mathcal{I}}$ and $M_{i}\in F(X_{i})^{c}$ that lifts $M$.
@@
The first assertion follows from [GR17, Cor. 7.2.7], where the colimit on the right is taken in $\textup{Cat}_{\infty}$.
Note that the filtered colimit of idempotent complete $\infty$-categories is idempotent complete [Lur09, Cor. 4.4.5.21]. Thus, any compact object comes from an object of the colimit on the right hand side (cf. [Lur09, Lem. 5.4.2.4] and [Lur21, 02LG]). ∎
@@
If $f\colon X\rightarrow Y$ is a morphism in $\operatorname{Fun}({\mathcal{I}},{\mathcal{E}}^{\operatorname{op}})$ such that each $f_{i}$ admits a left adjoint for all $i\in{\mathcal{I}}$ and each square is adjointable (in the sense of Proposition 3.7), then $f^{*}\coloneqq F_{{\mathcal{I}}}(f)$ admits a left adjoint.
@@
By Lemma 2.10 (1), we see that $f_{*}\colon\lim_{{\mathcal{I}}}F_{*}\circ X\rightarrow\lim_{{\mathcal{I}}}F_{*}\circ Y$ admits a left adjoint $f^{*}$, that has to be compatible with the identification of $\lim_{{\mathcal{I}}}F_{*}$ and $F_{{\mathcal{I}}}$. We claim that the following diagram is commutative (up to homotopy)
[…]
for all $i\in{\mathcal{I}}$. This claim implies that $f^{*}$ commutes with limits and therefore admits a left adjoint $f_{\sharp}$.